Given that Tan y = 2/5, find 4/5Cos y - 3/4Sin y
Answer by lwsshak3(11628)

Given that Tan y = 2/5, find 4/5Cos y - 3/4Sin y
tany=2/5=opposite side/adjacent side
hypotenuse=√(2^2+5^2)=√(4+25)=√29
cosy=5/√29
siny=2/√29
4/5Cos y-3/4Sin y
=(4/5)(5/√29)-(3/4)(2/√29)
=(4√29)–(3/2√29)
=(8-3)/2√29
=5/(2√29)